$SPAD/input schaum26.input

Timothy Daly December 3, 2016

Contents

2

1 [?]:14.525 ∫ lnx dx

∫ lnx = x ln x − x

— * —

)set break resume
)spool schaum26.output
)set message test on
)set message auto off
)clear all
--S 1 of 43
aa:=integrate(log(x),x)
--R
--R
--R   (1)  x log(x) - x
--R                                         Type: Union(Expression(Integer),...)
--E
--S 2 of 43
bb:=x*log(x)-x
--R
--R   (2)  x log(x) - x
--R                                                    Type: Expression(Integer)
--E
--S 3 of 43      14:525 Schaums and Axiom agree
cc:=aa-bb
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E

———-

2 [?]:14.526 ∫ x ln x dx

∫ x ln x = \frac{x^{2}}{2} (ln x − \frac{1}{2})

— * —

)clear all
--S 4 of 43
aa:=integrate(x*log(x),x)

--R
--R
--R          2          2
--R        2x log(x) - x
--R   (1)  --------------
--R               4
--R                                         Type: Union(Expression(Integer),...)
--E
--S 5 of 43
bb:=x^2/2*(log(x)-1/2)
--R
--R          2          2
--R        2x log(x) - x
--R   (2)  --------------
--R               4
--R                                                    Type: Expression(Integer)
--E
--S 6 of 43      14:526 Schaums and Axiom agree
cc:=aa-bb
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E

———-

3 [?]:14.527 ∫ x^{m} ln x dx

∫ x^{m} ln x = \frac{x^{m+1}}{m + 1} (ln x − \frac{1}{m + 1})

— * —

)clear all
--S 7 of 43
aa:=integrate(x^m*log(x),x)
--R
--R
--R                               m log(x)
--R        ((m + 1)x log(x) - x)%e
--R   (1)  -------------------------------
--R                   2
--R                  m  + 2m + 1
--R                                         Type: Union(Expression(Integer),...)
--E

--S 8 of 43
bb:=x^(m+1)/(m+1)*(log(x)-1/(m+1))
--R
--R                            m + 1
--R        ((m + 1)log(x) - 1)x
--R   (2)  -------------------------
--R                2
--R               m  + 2m + 1
--R                                                    Type: Expression(Integer)
--E
--S 9 of 43
cc:=aa-bb
--R
--R                               m log(x)                         m + 1
--R        ((m + 1)x log(x) - x)%e         + ((- m - 1)log(x) + 1)x
--R   (3)  -------------------------------------------------------------
--R                                  2
--R                                 m  + 2m + 1
--R                                                    Type: Expression(Integer)
--E
--S 10 of 43
explog:=rule(%e^(n*log(x)) == x^n)
--R
--R          n log(x)     n
--R   (4)  %e         == x
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 11 of 43
dd:=explog cc
--R
--R                              m + 1                         m
--R        ((- m - 1)log(x) + 1)x      + ((m + 1)x log(x) - x)x
--R   (5)  -----------------------------------------------------
--R                              2
--R                             m  + 2m + 1
--R                                                    Type: Expression(Integer)
--E
--S 12 of 43     14:527 Schaums and Axiom agree
ee:=complexNormalize dd
--R
--R   (6)  0
--R                                                    Type: Expression(Integer)
--E

———-

4 [?]:14.528 ∫ \frac{ln x}{x} dx

∫ \frac{ln x}{x} = \frac{1}{2} ln^{2} x

— * —

)clear all
--S 13 of 43
aa:=integrate(log(x)/x,x)
--R
--R
--R              2
--R        log(x)
--R   (1)  -------
--R           2
--R                                         Type: Union(Expression(Integer),...)
--E
--S 14 of 43
bb:=1/2*log(x)^2
--R
--R              2
--R        log(x)
--R   (2)  -------
--R           2
--R                                                    Type: Expression(Integer)
--E
--S 15 of 43     14:528 Schaums and Axiom agree
cc:=aa-bb
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E

———-

5 [?]:14.529 ∫ \frac{ln x}{x^{2}} dx

∫ \frac{ln x}{x^{2}} = −\frac{ln x}{x} − \frac{1}{x}

— * —

)clear all
--S 16 of 43
aa:=integrate(log(x)/x^2,x)
--R
--R
--R        - log(x) - 1
--R   (1)  ------------
--R              x
--R                                         Type: Union(Expression(Integer),...)
--E
--S 17 of 43
bb:=-log(x)/x-1/x
--R
--R        - log(x) - 1
--R   (2)  ------------
--R              x
--R                                                    Type: Expression(Integer)
--E
--S 18 of 43     14:529 Schaums and Axiom agree
cc:=aa-bb
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E

———-

6 [?]:14.530 ∫ ln^{2} x dx

∫ ln^{2} x = x ln^{2} x − 2x ln x + 2x

— * —

)clear all
--S 19 of 43
aa:=integrate(log(x)^2,x)
--R
--R
--R                2
--R   (1)  x log(x)  - 2x log(x) + 2x
--R                                         Type: Union(Expression(Integer),...)
--E
--S 20 of 43

bb:=x*log(x)^2-2*x*log(x)+2*x
--R
--R                2
--R   (2)  x log(x)  - 2x log(x) + 2x
--R                                                    Type: Expression(Integer)
--E
--S 21 of 43     14:530 Schaums and Axiom agree
cc:=aa-bb
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E

———-

7 [?]:14.531 ∫ \frac{ln^{n} x}{x} dx

∫ \frac{ln^{n} x}{x} = \frac{ln^{n+1}x}{n + 1}

— * —

)clear all
--S 22 of 43
aa:=integrate(log(x)^n/x,x)
--R
--R
--R                n log(log(x))
--R        log(x)%e
--R   (1)  ---------------------
--R                n + 1
--R                                         Type: Union(Expression(Integer),...)
--E
--S 23 of 43
bb:=log(x)^(n+1)/(n+1)
--R
--R              n + 1
--R        log(x)
--R   (2)  -----------
--R           n + 1
--R                                                    Type: Expression(Integer)
--E
--S 24 of 43
cc:=aa-bb

--R
--R                n log(log(x))         n + 1
--R        log(x)%e              - log(x)
--R   (3)  -----------------------------------
--R                       n + 1
--R                                                    Type: Expression(Integer)
--E
--S 25 of 43
explog:=rule(%e^(n*log(x)) == x^n)
--R
--R          n log(x)     n
--R   (4)  %e         == x
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 26 of 43
dd:=explog cc
--R
--R                n + 1               n
--R        - log(x)      + log(x)log(x)
--R   (5)  -----------------------------
--R                    n + 1
--R                                                    Type: Expression(Integer)
--E
--S 27 of 43     14:531 Schaums and Axiom agree
ee:=complexNormalize dd
--R
--R   (6)  0
--R                                                    Type: Expression(Integer)
--E

———-

8 [?]:14.532 ∫ \frac{dx}{x ln x}

∫ \frac{1}{x ln x} = ln(ln x)

— * —

)clear all
--S 28 of 43
aa:=integrate(1/(x*log(x)),x)
--R
--R

--R   (1)  log(log(x))
--R                                         Type: Union(Expression(Integer),...)
--E
--S 29 of 43
bb:=log(log(x))
--R
--R   (2)  log(log(x))
--R                                                    Type: Expression(Integer)
--E
--S 30 of 43     14:532 Schaums and Axiom agree
cc:=aa-bb
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E

———-

9 [?]:14.533 ∫ \frac{dx}{ln x}

∫ \frac{1}{ln x} = ln(ln x) + ln x + \frac{ln^{2} x}{2 · 2!} + \frac{ln^{3} x}{3 · 3!} + ⋯

— * —

)clear all
--S 31 of 43     14:533 Schaums and Axiom agree by definition
aa:=integrate(1/log(x),x)
--R
--R
--R   (1)  li(x)
--R                                         Type: Union(Expression(Integer),...)
--E

———-

10 [?]:14.534 ∫ \frac{x^{m}}{ln x} dx

∫ \frac{x^{m}}{ln x} = ln(ln x) + (m + 1) ln x + \frac{(m + 1)^{2} ln^{2} x}{2 · 2!} + \frac{(m + 1)^{3} ln^{3} x}{3 · 3!} + ⋯

— * —

--S 32 of 43     14:534 Axiom cannot compute this integral
aa:=integrate(x^m/log(x),x)
--R
--R
--R           x     m
--I         ++    %I
--I   (1)   |   ------- d%I
--I        ++   log(%I)
--R                                         Type: Union(Expression(Integer),...)
--E

———-

11 [?]:14.535 ∫ ln^{n} x dx

∫ ln^{n} x = x ln^{n} x − n ∫ ln^{n−1} x

— * —

)clear all
--S 33 of 43     14:535 Axiom cannot compute this integral
aa:=integrate(log(x)^n,x)
--R
--R
--R           x
--R         ++         n
--I   (1)   |   log(%I) d%I
--R        ++
--R                                         Type: Union(Expression(Integer),...)
--E

———-

12 [?]:14.536 ∫ x^{m} ln^{n} x dx

∫ x^{m} ln^{n} x = \frac{x^{m+1} ln^{n} x}{m + 1} − \frac{n}{m + 1} ∫ x^{m} ln^{n−1} x

— * —

)clear all
--S 34 of 43     14:536 Axiom cannot compute this integral

aa:=integrate(x^m*log(x)^n,x)
--R
--R
--R           x
--R         ++    m       n
--I   (1)   |   %I log(%I) d%I
--R        ++
--R                                         Type: Union(Expression(Integer),...)
--E

———-

13 [?]:14.537 ∫ ln (x^{2} + a^{2}) dx

∫ ln (x^{2} + a^{2}) = x ln(x^{2} + a^{2}) − 2x + 2a tan^{−1} \frac{x}{a}

— * —

)clear all
--S 35 of 43
aa:=integrate(log(x^2+a^2),x)
--R
--R
--R               2    2            x
--R   (1)  x log(x  + a ) + 2a atan(-) - 2x
--R                                 a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 36 of 43
bb:=x*log(x^2+a^2)-2*x+2*a*atan(x/a)
--R
--R               2    2            x
--R   (2)  x log(x  + a ) + 2a atan(-) - 2x
--R                                 a
--R                                                    Type: Expression(Integer)
--E
--S 37 of 43     14:537 Schaums and Axiom agree
cc:=aa-bb
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E

———-

14 [?]:14.538 ∫ ln(x^{2} − a^{2}) dx

∫ ln(x^{2} − a^{2}) = x ln(x^{2} − a^{2}) − 2x + a ln (\frac{x + a}{x − a})

— * —

)clear all
--S 38 of 43
aa:=integrate(log(x^2-a^2),x)
--R
--R
--R               2    2
--R   (1)  x log(x  - a ) + a log(x + a) - a log(x - a) - 2x
--R                                         Type: Union(Expression(Integer),...)
--E
--S 39 of 43
bb:=x*log(x^2-a^2)-2*x+a*log((x+a)/(x-a))
--R
--R               2    2          x + a
--R   (2)  x log(x  - a ) + a log(-----) - 2x
--R                               x - a
--R                                                    Type: Expression(Integer)
--E
--S 40 of 43
cc:=aa-bb
--R
--R                                            x + a
--R   (3)  a log(x + a) - a log(x - a) - a log(-----)
--R                                            x - a
--R                                                    Type: Expression(Integer)
--E
--S 41 of 43     14:538 Schaums and Axiom agree
dd:=expandLog cc
--R
--R   (4)  0
--R                                                    Type: Expression(Integer)
--E

———-

15 [?]:14.539 ∫ x^{m} ln(x^{2} ± a^{2}) dx

∫ x^{m} ln(x^{2} ± a^{2}) = \frac{x^{m−1} ln(x^{2} ± a^{2})}{m + 1} − \frac{2}{m + 1} ∫ \frac{x^{m+2}}{x^{2} ± a^{2}}

— * —

)clear all
--S 42 of 43
aa:=integrate(x^m*log(x^2+a^2),x)
--R
--R
--R           x
--R         ++       2     2   m
--I   (1)   |   log(a  + %I )%I d%I
--R        ++
--R                                         Type: Union(Expression(Integer),...)
--E
)clear all
--S 43 of 43     14:539 Axiom cannot compute this integral
aa:=integrate(x^m*log(x^2-a^2),x)
--R
--R
--R           x
--R         ++         2     2   m
--I   (1)   |   log(- a  + %I )%I d%I
--R        ++
--R                                         Type: Union(Expression(Integer),...)
--E
)spool
)lisp (bye)

———-

References

[1] Spiegel, Murray R. Mathematical Handbook of Formulas and Tables Schaum’s Outline Series McGraw-Hill 1968 p86